Ariyone claims that 3,6 is the point of intersection of the lines y= 4x-2 and y=1/2x+5. is he correct. How do you know?
Found 3 solutions by scott8148, nerdybill, Alan3354:
Answer by scott8148(6628)

Ariyone is "misguided" at best

substituting the point (3,6) into both equations, shows that the point is not on either line

Answer by nerdybill(7384)

Ariyone claims that 3,6 is the point of intersection of the lines y= 4x-2 and y=1/2x+5. is he correct. How do you know?
.
The claim is, given two lines:
y= 4x-2
y=1/2x+5
.
It intersects at (x,y)=(3,6)
.
plugging the point into:
y= 4x-2
6= 4(3)-2
6= 12-2
6= 10 (false)
.
plugging the point into:
y=1/2x+5
6=1/2(3)+5
6=3/2+5
6=3/2+10/2
6=13/2 (false)
.
So, NO, he is incorrect.
The intersection can be found by setting both equations equal to each other:
4x-2 =1/2x+5
8x-4 = x+10
7x-4 = 10
7x = 14
x = 2
.
plug it back into either equation to find y:
y = 4x-2
y = 4(2)-2
y = 8-2
y = 6
.
Intersecting point is at (2,6)

Answer by Alan3354(69443)

Ariyone claims that 3,6 is the point of intersection of the lines y= 4x-2 and y=1/2x+5. is he correct. How do you know?
0 solutions
---------------
Sub 3 for x and 6 for y into the equations.
y = 4x-2
6 =? 4*3 - 2
6<>10, so it's not on the line.
There's no need to check the other one, since he's been proven wrong.